You plan to invest $19,000 in two funds paying 3% and 4.5
simple interest. (There is more risk in the 4.5 fund) Your goal is to obtain a total annual interest income of $700 from the investments. What is the smallest amount you can invest in the 4.5fund and still meet your objective?

Answer by richwmiller(17219)

x+y=19000,
0.03*x+0.045*y=700
x=19000-y
0.03*(19000-y)+0.045*y=700
570-0.03y+0.045*y=700
0.015*y=130
y=8666.67 at 4.5%
x=19000-y
x=10333.33 at 3%
check
0.03*10333.33+0.045*8666.67=700
310.0+390.0=700
700.0=700
ok